# VisionReward: Fine-Grained Multi-Dimensional Human Preference Learning for Image and Video Generation

### v2

@@ 3.1 VisionReward for Text-to-Vision Evaluation
Dataset & Training Setting. After balanced sampling, we obtain 40,743 images and corresponding training data containing 97,680 judgment questions from images, leaving 6,910 images for subsequent validation and test. For videos, we obtain 28,605 videos and corresponding training data containing 89,473 judgment questions, with 3,080 videos reserved.
@@
To fine-tune CogVLM2 [18], we set a batch size of 64, a learning rate of 1e-6, and train for 1,500 steps. For CogVLM-Video, we set a batch size of 64, a learning rate of 4e-6, and train for 1,500 steps.
@@
To learn linear weights for preference prediction, we sample human preference pairs and perform logistic regression. For images, We sample 44k pairs (24k from HPDv2 [47] and 20k from ImageRewardDB [49]); and for videos, we sample prompts from VidProM [44] and generate videos (using CogVideoX [50], VideoCrafter2 [4] and OpenSora [54]), getting 1,795 annotated video pairs with preference.
@@
To establish a comprehensive evaluation benchmark for both image and video generation, we construct MonetBench, which contains separate test sets for images and videos, each consisting of 1,000 prompts.  Tab. 1 shows content and challenge categories for MonetBench. More details are introduced in Appendix  Sec. 11.
@@
Main Results: Preference Accuracy. Preference accuracy means the probability that a reward model has the same judgment as humans about which image is better. We use MonetBench to construct our test set for human preference, using SDXL [31] to generate images and CogVideoX [50] / VideoCrafter2 [4] / OpenSora [54] to generate videos, resulting in 500 pairs for image and 1,000 pairs for video. We employ annotators to assess the generated images using a preference rating scale from 1 to 5 (with 3 indicating no preference). The average preference score is used as the final preference label. We also take HPDv2 [47] and GenAI-Bench [20] as test set.
@@
Tab. 2 shows that VisionReward obtains state-of-the-state results in multiple datasets. Notably, in video evaluation, image reward models demonstrate competitive performance when the video duration is within 2 seconds (GenAI-Bench). However, when the video duration reaches 6 seconds (MonetBench), only VisionReward is capable of accurately predicting human preference, being twice (22.1% over random) as high as the best (12.5% over random) among other methods. This indicates that dynamic information in longer videos poses a challenge for RMs, while VisionReward can effectively address this issue after fine-grained visual learning.
Main Results: Accuracy and Consistency on Judgment. To evaluate the effectiveness of judgment learning, we construct a visual quality QA set to assess the visual assessment capabilities of VisionReward compared to other VLMs. We collect 1,364 test cases for images involving 14 types of questions across 4 dimensions, and additionally 1,308 cases covering 8 types of questions across 4 dimensions related to dynamic content in videos. To ensure the generality of these questions, we combine adjacent degrees in the checklists under each sub-dimension, enhancing distinctiveness and minimizing incidental subjectivity.  Tab. 3 demonstrates VisionReward’s superiority in judging visual quality over existing generalized multimodal LLMs, which remains a challenge for generalized multimodal LLMs.
As questions corresponding to each sub-dimension assess varying degrees of a particular factor, it’s important to measure consistency of VisionReward across multiple questions of the same sub-dimension. Consistency measures the likelihood that the model provides consistent responses across a series of judgments concerning this factor.  Tab. 4 shows that VisionReward has high consistency (more than 97 %) in most (7 of 8) dimensions.
Ablation Study: Impact of Training Set Size. We conduct experiments with varying sizes of the training set to investigate its influence on regression performance. As demonstrated in  Tab. 5, the accuracy improves monotonically as the training set size increases up to 4,000 samples, beyond which the accuracy stabilizes.
@@
Ablation Study: Scalability of Question Scale.  Tab. 6 shows that accuracy of preference prediction exhibits significant improvement as the question scale increases. Scalability validates the effects of decomposing preferences via multi-scale questioning. We do more analysis such as question pruning in Appendix  Sec. 8.
Size 200 500 1k 2k 4k 8k
Acc. 77.6 80.3 80.6 80.9 81.3 81.3
Table 5: Accuracy on HPDv2-test for different sizes of train set.
#question 2 4 8 16 32 all
Acc. 59.3 68.6 72.5 75.8 78.9 81.7
Table 6: Accuracy on HPDv2-test for different question scales.
3.2 VisionReward for Text-to-Vision Optimization
Figure 4: Human evaluation results of text-to-image DPO.
Figure 5: Human evaluation results of text-to-video DPO.
@@ 3.2 VisionReward for Text-to-Vision Optimization
To evaluate the efficacy of VisionReward in preference optimization for visual generative systems, we conducted a series of comparative experiments against current state-of-the-art reward models and established preference datasets. The empirical results presented in  Fig. 4 and  Fig. 5 demonstrate VisionReward’s superior performance in both text-to-image and text-to-video optimization, achieving statistically significant improvements in human preference metrics over competing approaches.
@@
This section focuses on preference optimization for text-to-video using different reward models. Details of text-to-image are provided in Appendix  Sec. 9.2 due to page limits.
@@
Dataset & Training Settings. For our backbone model, we select CogVideoX-2B [^1] https://huggingface.co/THUDM/CogVideoX-2b. The training prompts are sampled from VidProM [44]. To adapt these prompts for video generation, we have optimized them following guidelines from CogVideoX [50], which results in roughly 22,000 samples. We generate 4 videos for each prompt, and use VisionReward to score these videos and apply the MPO strategy to select approximately 9,400 effective preference pairs. In all our experiments, we maintain a batch size of 32, a learning rate of 5e-6, and employ 100 warmup steps followed by linear decay. We set the DPO parameter $\beta$ to 500. The MPO training process spans around 500 steps, equivalent to about 2 epochs. During training, we save a checkpoint every 40 steps and use a validation set split from the training set to pick the checkpoint with the highest reward.
@@
Evaluation Settings. To comprehensively assess the MPO models, we have conducted both automatic and human evaluations. The automatic evaluation is conducted across various benchmarks, including VBench [19] and our Video-MonetBench. For VBench, we focus on commonly reported key metrics, including Human Action, Scene, Multiple Objects, and Appearance Style (Appear. Style). For human evaluation, annotators are asked to watch two videos and select the better one, with an option to indicate a tie if neither stands out. This manual assessment also uses the tasks in MonetBench. In all these experiments, we utilize prompt optimization recommended in CogVideoX. Our baseline comparisons include the original CogVideoX-2B and DPO with VideoScore.
[…]
Style
@@
Experimental Results. The main results are shown in  Tab. 7 and  Tab. 8. When compared to the original CogVideoX-2B, optimization with VisionReward significantly enhances model performance across these benchmarks. In contrast, optimization with VideoScore tends to degrade performance. The empirical evidence substantiates VisionReward’s advanced capacity for multi-dimensional optimization. (Case study in Appendix  Sec. 9.3.)
3.3 Effects of Multi-Dimensional Optimization
Ablation Study: Different Strategy for MPO. In using the MPO strategy, we define the way in which $R^{i}$ dominates $R^{j}$ (given two images $x^{i}$ and $x^{j}$) to select pairs. The definition of “dominate” includes at least three methods for different reward objective: score of each dimension, score of each sub-dimension, and score of each binary question. To investigate the impact of different definitions of “dominate”, we conduct experiments based on CogVideoX-5B. Specifically, we employ three different strategies, setting the threshold of the total score to 0.6, 0.5, and 0.4 respectively, ensuring that the number of pairs obtained through all three strategies is 5k. We use a batch size of 64, a learning rate of 2e-6, the DPO parameter $\beta$ of 500, and the training steps of 300. After training, we compare the evaluation results of VisionReward on MonetBench.  Tab. 9 shows that using the score for each dimension as objective yields the best results.
Methods Baseline Dimension Sub-dimension Question
Reward 4.303 4.573 4.539 4.514
@@ 3.3 Effects of Multi-Dimensional Optimization
Table 10: Ablation Study for MPO strategy. Scores are given by VisionReward on MonetBench.
Ablation Study: Multi-Dimensional or Single Score. To comprehensively illustrate how MPO addresses the factor optimization bias inherent in DPO, we extend the previously described experimental setup by utilizing the total score of VisionReward to select pairs, setting the threshold at 0.8 to ensure consistency in numbers. VisionReward is employed to evaluate scores across various dimensions, with the detailed results presented in  Tab. 10. Through the implementation of the MPO strategy, CogVideoX is optimized in such a manner that it avoids the degradation of certain factors (e.g., alignment), thereby achieving improved trade-offs, such as maintaining good preservation while avoiding excessively slow dynamic changes. The empirical evidence further substantiates VisionReward’s capacity for algorithm-agnostic preference alignment, as evidenced by comparative testing with other approaches like MaPO [16].
For efficiency discussion, the experimental results in  Tab. 11 reveal the comparative effectiveness of the MPO and DPO methods. These results suggest that MPO outperforms the DPO approach in terms of both efficiency and effectiveness.
These findings highlight promising directions for future research in developing novel optimization algorithms and adapting VisionReward for multi-dimensional optimization and preference balancing.
Method Threshold #Sample Overall
DPO 0.6 6473 4.442
DPO 0.8 5270 4.515
MPO 0.6 5028 4.573
Table 11: More comparison of MPO and DPO on MonetBench.
4 Related Work
Evaluation for Text-to-image and Text-to-video. The development of multimodal generative models is advancing rapidly, including text-to-image [9, 33, 35, 34, 2, 31] and text-to-video [17, 15, 42, 54, 4, 50] models that generate high-quality images or videos based on given textual input. For text-to-image, early evaluations heavily relied on metrics such as FID [13] and CLIP [32] Score. Recent work [49] has pointed out that these metrics may not align well with human preferences. Many metrics based on reward models learn from human preferences [49, 22, 47], resulting in a single score, which can lead to biases. With the advancement of Vision-Language Models (VLMs) [28, 45, 24, 18, 1, 41], some studies [53, 27] have attempted to directly use VLMs to assess images, considering aspects like alignment and quality. Nevertheless, we find that VLMs have a gap in their ability to accurately assess vision quality. VisionReward proposes a fine-grained, multi-dimensional evaluation framework, using specialized judgment instructions to achieve more accurate and interpretable.
RLHF for Text-to-image and Text-to-video. Reinforcement Learning from Human Feedback (RLHF) [40, 29, 30] refers to optimizing models with reinforcement learning based on human feedback, which has been proven to significantly enhance language models. For visual generation tasks, several works have explored RLHF, optimizing from the gradient [49, 48] or using a policy-based RL approach [3, 10, 7]. All these methods require a reward model (RM) to provide feedback for online learning. Diffusion-DPO [43] has proposed to optimize the diffusion model directly using human-labeled preference data. However, most RLHF methods face the issue of over-optimization. VisionReward highlights that these issues are because of excessive optimization of some factors and degradation of others. By employing a Pareto-optimality approach from a multi-objective perspective, VisionReward achieves robust RLHF.
@@ 6 Acknowledgments
The authors would like to thank Zhipu AI for sponsoring the computation resources used in this work.
@@
7 Details of Annotation and Dataset
7.1 Details of Annotation Pipeline
Details of Data Preparation. To ensure the diversity of our annotation, we collect our annotated data from various sources, as presented in  Tab. 12. For images, we select 3 popular datasets: ImageRewardDB [49], HPDv2 [47], and Pick-a-Pic [22]. We sample 16k images from each dataset for annotation (with the first two sampling 4k prompts each associated with 4 images, and Pick-a-Pic sampling 8k prompts each associated with 2 images), totaling 48k annotated images. For videos, we conduct random sampling from the VidProM dataset and use ChatGLM [11] for data cleaning (prompt Cf.  Tab. 13), leading to 10k prompts. Then we use CogVideoX [50], VideoCrafter2 [4] and OpenSora [54] to generate 30k videos, sample from Panda-70M [5] to get 3k real videos, leading to 33k videos for annotation.
@@ 7.1 Details of Annotation Pipeline
Image ImageRewardDB [49] 16K 1M
Pick-a-Pic [22] 16K 1M
HPDv2 [47] 16K 1M
Video CogVideoX [50] 10K 0.6M
Open-Sora [54] 10K 0.6M
VideoCrafter2 [4] 10K 0.6M
Panda-70M [5] 3K 0.2M
@@
(a) Text-to-image
(b) Text-to-video
Figure 6: Annotation statistics of different sub-dimensions.
Annotation Design. In the annotation process, we design a comprehensive dimensional system and divide each dimension into several sub-dimensions.  Tab. 14 illustrates our preference dimensions and the checklist count. We identify 5 dimensions for text-to-image generation and expand to 9 dimensions for text-to-video. For each sub-dimension, we set options that vary gradually in degree. Annotators are required to label each sub-dimension for every image. In our annotation process, we design multiple-choice questions for each dimension, allowing annotators to select the option that best fits the image. Additionally, we further decompose these options into several binary (yes/no) questions, ultimately forming a checklist for evaluating each image or video. This checklist is then utilized for subsequent evaluation systems and reward model training.
Annotation Management. To avoid annotation bias of annotators, our approach includes:
  • Professional management. Cooperated with a specialized company, we strictly conduct trial annotation training for the annotators, screen qualified annotators, and conduct secondary quality inspection of the annotation results.
  • Standard document. Our annotation document gives clear definitions and provides more than 10 examples for each judgment, to align the standard among annotators.
  • Consistency test. Consistency of annotators on checklist results in 89.29% (images) and 89.33% (videos).
USER Here is the prompt you have received: [[PROMPT]]
Table 13: Prompt template and example for prompt cleaning.
@@ 8 More Details and Results of VisionReward
Visual Judgment Process. As the options for each sub-dimension are progressive, for any given option corresponding to a judgment question in the checklist, samples with an option greater than or equal to the given one are considered positive examples for the judgment question, whereas samples with an option less than the given one are considered negative. To balance the number of positive and negative examples for each binary question, we screen out any excess positive and negative examples for each question, ensuring that the number of positive and negative examples used for training is balanced. For alignment, we use different methods on images and videos. For images, we use VQAScore [27] as an alignment judgment. For videos, we train five levels of judgment for VisionReward.
@@
4:     $_{\text{old}}\leftarrow\textbf{w}$ w
@@
10:     Mask negative weights $\mathbf{w}\leftarrow\mathbf{w}\odot(\mathbf{w}>0)$
11:     $\text{diff}\leftarrow||\textbf{w}-\textbf{w}_{\text{old}}||$
11:  Trained weights $\mathbf{w}$
Ablation Study: Efficiency of VisionReward via Weight Masking. We explore whether selective question pruning through coefficient thresholding could maintain model efficacy. As demonstrated in  Tab. 15, coefficient-guided masking strategy (Acc. w/ mask) achieves remarkable efficiency gains with comparable accuracy.
#question 2 4 8 16 32 all
Acc. w/ mask 71.1 78.9 80.7 81.4 81.6 81.7
Table 15: Accuracy on HPDv2-test for different question scales. “Acc. w/ mask” means selecting questions by coefficient magnitude after full-set regression.

### v3

@@ 4.1 VisionReward for Text-to-Vision Evaluation
Dataset & Training Setting. After balanced sampling, we obtain 40,743 images and corresponding 97,680 judgment questions for training, leaving 6,910 images for subsequent validation and test. For videos, we obtain 28,605 videos and corresponding 89,473 judgment questions for training, with 3,080 videos reserved.
@@
To fine-tune CogVLM2 (hong2024cogvlm2), we set a batch size of 64, a learning rate of 1e-6, and train for 1,500 steps. For CogVLM-Video, we set a batch size of 64, a learning rate of 4e-6, and train for 1,500 steps.
@@
To learn linear weights for preference prediction, we sample human preference pairs and perform logistic regression. For images, We sample 44k pairs (24k from HPDv2 (wu2023human) and 20k from ImageRewardDB (xu2023imagereward)); and for videos, we sample prompts from VidProM (wang2024vidprom) and generate videos (using CogVideoX (yang2024cogvideox), VideoCrafter2 (chen2024videocrafter2) and OpenSora (opensora)), getting 1,795 annotated video pairs with preference.
@@
To establish a comprehensive evaluation benchmark for both image and video generation, we construct MonetBench, which contains separate test sets for images and videos, each consisting of 1,000 prompts. More details are introduced in Appendix Section 13.
Figure 5: The accuracy of VisionReward improves as the number of binary questions increases. After masking weights from full regression, VisionReward maintains high performance.
Figure 6: Human evaluation results of DPO using different reward models or preference datasets. We require five annotators to comprehensively evaluate two samples and select the better one. VisionReward achieve the best performance.
@@
Main Results: Preference Accuracy. Preference accuracy means the probability that a reward model has the same judgment as humans about which image is better. We use MonetBench to construct our test set for human preference, using SDXL (podell2023sdxl) to generate images and CogVideoX (yang2024cogvideox) / VideoCrafter2 (chen2024videocrafter2) / OpenSora (opensora) to generate videos, resulting in 500 pairs for image and 1,000 pairs for video. We employ annotators to assess the generated images using a preference rating scale from 1 to 5 (with 3 indicating no preference). The average preference score is used as the final preference label. We also take HPDv2 (wu2023human) and GenAI-Bench (jiang2024genai) as test set.
@@
Table 2 shows that VisionReward obtains state-of-the-state results in multiple datasets. Notably, in video evaluation, image reward models demonstrate competitive performance when the video duration is within 2 seconds (GenAI-Bench). However, when the video duration reaches 6 seconds (MonetBench), only VisionReward is capable of accurately predicting human preference, being twice (22.1% over random) as high as the best (12.5% over random) among other methods. This indicates that dynamic information in longer videos poses a challenge for RMs, while VisionReward can effectively address this issue after fine-grained visual learning.
@@
Ablation Study: Scalability of Question Scale.  Fig. 5 shows that accuracy of preference prediction exhibits significant improvement as the question scale increases, and masking minimal weights maintains performance (Details in Appendix  Section 8). Scalability validates the effects of decomposing preferences via fine-grained questions.
We do more analysis of VisionReward in Appendix  Section 8 and analysis of fine-grained results in  Section 11.
4.2 VisionReward for Preference Optimization
@@
To evaluate the efficacy of VisionReward in preference optimization for visual generative systems, we conduct a series of comparative experiments against current state-of-the-art reward models and established preference datasets. We use SDXL as text-to-image base model and CogVideoX as text-to-video base model. The empirical results presented in  Fig. 6 demonstrate VisionReward’s superior performance, achieving statistically significant improvements in human preference metrics over competing approaches.
@@
This section focuses on our experiments on preference optimization for text-to-video using different reward models. Details of experiments for text-to-image are provided in Appendix  Section 9.2 due to page limits.
@@
Dataset & Training Settings. For our backbone model, we select CogVideoX-2B. The training prompts are sampled from VidProM (wang2024vidprom) (details in Appendix Section 9.1). To adapt these prompts for video generation, we have optimized them following guidelines from CogVideoX (yang2024cogvideox), which results in roughly 22,000 samples. We generate 4 videos for each prompt, and use VisionReward to score these videos and apply the MPO strategy to select approximately 9,400 effective preference pairs. In all our experiments, we maintain a batch size of 32, a learning rate of 5e-6, and employ 100 warmup steps followed by linear decay. We set the DPO parameter $\beta$ to 500. The MPO training process spans around 500 steps, equivalent to about 2 epochs. During training, we save a checkpoint every 40 steps and use a validation set split from the training set to pick the checkpoint with the highest reward.
@@
Evaluation Settings. To comprehensively assess the MPO models, we have conducted both automatic and human evaluations. The automatic evaluation is conducted across various benchmarks, including VBench (huang2024vbench) and our Video-MonetBench. For VBench, we focus on commonly reported key metrics, including Human Action, Scene, Multiple Objects, and Appearance Style. In all these experiments, we utilize prompt optimization recommended in CogVideoX. Our baseline comparisons include the original CogVideoX-2B and DPO with VideoScore.
Methods    ActionHuman Scene    ObjectsMultiple    StyleAppear.
@@
Experimental Results. The main results are shown in  Table 3 and  Table 4. When compared to the original CogVideoX-2B, optimization with VisionReward significantly enhances model performance across these benchmarks. In contrast, optimization with VideoScore tends to degrade performance. The empirical evidence substantiates VisionReward’s advanced capacity for multi-dimensional optimization. (Case study in Appendix Section 9.3.)
4.3 Ablation Study of MPO
To comprehensively illustrate how MPO addresses the factor optimization bias inherent in DPO, We conduct experiments based on CogVideoX-5B. We set the threshold of the total score to 0.8 for DPO and 0.6 for MPO, ensuring that the number of pairs obtained through all three strategies is 5k. We use a batch size of 64, a learning rate of 2e-6, the DPO parameter $\beta$ of 500, and the training steps of 300. VisionReward is employed to evaluate scores across various dimensions, with the detailed results presented in  Table 5. Through the implementation of the MPO strategy, CogVideoX is optimized in such a manner that it avoids the degradation of certain factors (e.g., alignment), thereby achieving improved trade-offs, such as maintaining good preservation while avoiding excessively slow dynamic changes. The empirical evidence further substantiates VisionReward’s capacity for algorithm-agnostic preference alignment, as evidenced by comparative testing with other approaches like MaPO (hong2024margin).
@@ 4.3 Ablation Study of MPO
Table 5: Ablation Study of MPO strategy. Scores are given by VisionReward on MonetBench.
For efficiency discussion, the experimental results in  Table 6 reveal the comparative effectiveness of the MPO and DPO methods. We analyze the pairs selected in perspective of $R^{i}$ dominating $R^{j}$” mentioned in  Section 3.3. These results suggest that MPO outperforms the DPO approach in terms of both efficiency and effectiveness. These findings highlight promising directions for future research in developing novel optimization algorithms and adapting VisionReward for multi-dimensional optimization.
Method #Dom. #Not-Dom. Reward
Original - - 4.303
DPO w/o MPO 3814 1456 4.515 (+0.212)
DPO w/ MPO 5028 0 4.573 (+0.270)
$\Delta$ (vs w/o MPO) +31.8% -100% +27.4%
Table 6: Comparison of MPO and DPO on MonetBench. “#Dom.” means the number of pairs that match the rule of $R^{i}$ dominating $R^{j}$”, while “#Not-Dom.” pairs not match.
@@ 6 Acknowledgments
The authors would like to thank Z.AI for sponsoring the computation resources used in this work.
@@
7 More Details of Annotation
7.1 Details of Annotation Design
To ensure the diversity of our annotation, we collect our annotated data from various sources, as presented in  Table 7. The prompt and example for data cleaning is shown in  Table 9.  Table 8 illustrates our preference dimensions and the checklist count. We identify 5 dimensions for text-to-image generation and expand to 9 dimensions for text-to-video.
@@ 7.1 Details of Annotation Design
Image ImageRewardDB (xu2023imagereward) 16K 1M
Pick-a-Pic (kirstain2023pick) 16K 1M
HPDv2 (wu2023human) 16K 1M
Video CogVideoX (yang2024cogvideox) 10K 0.6M
Open-Sora (opensora) 10K 0.6M
VideoCrafter2 (chen2024videocrafter2) 10K 0.6M
Panda-70M (chen2024panda) 3K 0.2M
@@
(a) Text-to-image
(b) Text-to-video
Figure 7: Annotation statistics of different sub-dimensions.
USER Here is the prompt you have received: [[PROMPT]]
Table 9: Prompt template and example for prompt cleaning.
@@ 8 More Details and Results of VisionReward
Visual Judgment Process. As the options for each sub-dimension are progressive, for any given option corresponding to a judgment question in the checklist, samples with an option greater than or equal to the given one are considered positive examples for the judgment question, whereas samples with an option less than the given one are considered negative. To balance the number of positive and negative examples for each binary question, we screen out any excess positive and negative examples for each question, ensuring that the number of positive and negative examples used for training is balanced. For alignment, we use different methods on images and videos. For images, we use VQAScore (lin2025evaluating) as an alignment judgment. For videos, we train five levels of judgment for VisionReward.
Main Results: Accuracy on Judgment. To evaluate the effectiveness of judgment learning, we construct a visual quality QA set to assess the visual assessment capabilities of VisionReward compared to other VLMs. We collect 1,364 test cases for images involving 14 types of questions across 4 dimensions, and additionally 1,308 cases covering 8 types of questions across 4 dimensions related to dynamic content in videos. To ensure the generality of these questions, we combine adjacent degrees in the checklists under each sub-dimension, enhancing distinctiveness and minimizing incidental subjectivity.  Table 11 demonstrates VisionReward’s superiority in judging visual quality over existing generalized multimodal LLMs, which remains a challenge for generalized multimodal LLMs.
Composition Quality Fidelity Safety
97.9 98.2 98.3 99.1
Stability Dynamic Physics Preservation
97.4 99.9 88.2 99.8
Table 10: Consistency of VisionReward in each dimension.
Main Results: Consistency on Judgment. As questions corresponding to each sub-dimension assess varying degrees of a particular factor, it’s important to measure consistency of VisionReward across multiple questions of the same sub-dimension. Consistency measures the likelihood that the model provides consistent responses across a series of judgments concerning this factor.  Table 10 shows that VisionReward has high consistency (more than 97 %) in most (7 of 8) dimensions.
@@
4:  $\textbf{w}_{\text{old}}\leftarrow\textbf{w}$
@@
10:  end for
11:  Mask negative weights $\mathbf{w}\leftarrow\mathbf{w}\odot(\mathbf{w}>0)$
[…]
13:end while
13: Trained weights $\mathbf{w}$
Method Image Video
Composition Quality Fidelity Safety Stability Dynamic Physics Preservation
LLaVa 59.9 65.7 59.8 64.4 52.5 53.8 50.6 47.5
CogVLM2 (hong2024cogvlm2) 65.8 67.1 53.1 74.7 49.3 57.1 51.2 47.8
GPT-4o (achiam2023gpt) 73.1 62.7 61.9 70.1 57.9 69.1 62.4 58.8
Gemini (team2024gemini) 69.4 59.9 59.7 74.9 58.1 71.1 58.1 59.6
VisionReward (Ours) 78.8 81.1 80.9 83.9 64.8 75.4 68.1 72.0
Table 11: Accuracy of VisionReward and other vision-language models (VLMs) on vision quality questions constructed from our annotation. We test LLaVA-v1.5-7B (liu2023llava) for image and LLava-Next-Video-34B (li2024llava) for video.
Ablation Study: Impact of Training Set Size. We conduct experiments with varying sizes of the training set to investigate its influence on regression performance. As demonstrated in  Table 12, the accuracy improves monotonically as the training set size increases up to 4,000 samples, beyond which the accuracy stabilizes.
Size 200 500 1k 2k 4k 8k
Acc. 77.6 80.3 80.6 80.9 81.3 81.3
Table 12: Accuracy on HPDv2-test for different sizes of train set.